OpenStudy (anonymous):

the height of a triangle is 18 inches.find the length of the base if the area of the triangle is 63 square inches? please help me and show me steps

OpenStudy (tyteen4a03):

Remember the formula for area of triangles: \(\large{A = \frac{base * height}{2}}\).

OpenStudy (anonymous):

im stuck i plugged in everything and i dont know what to do now 63=B x 18/2

OpenStudy (anonymous):

the height of a triangle is 18 inches. find the length of the base if the area of a triangle is 63 square inches. Area of triangle = 1/2*base*height = 63 sqare inches. height = 18 inches Area = 1/2*base*18 = 63 base*18 = 63*2 base*18 = 126 base = 126/18 = 7 inches Answer: the length of base is 7 inches.

OpenStudy (anonymous):

thank you very helpful